OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 6 x – y = 3 Answer -2 -1 2 5

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

Simultaneous equations: Using elimination: 4x+2y=6 x-y=3 4x+2y=6 2x-2y=6 6x=12 x=2

OpenStudy (anonymous):